# Towards an Accessible, Noninvasive Micronutrient Status Assessment Method: A Comprehensive Review of Existing Techniques

### v1

@@ Abstract
Nutrients are critical to the functioning of the human body and their imbalance can result in detrimental health concerns. The majority of nutritional literature focuses on macronutrients, often ignoring the more critical nuances of micronutrient balance, which require more precise regulation. Currently, micronutrient status is routinely assessed via complex methods that are arduous for both the patient and the clinician. To address the global burden of micronutrient malnutrition, innovations in assessment must be accessible and noninvasive. In support of this task, this article synthesizes useful background information on micronutrients themselves, reviews the state of biofluid and physiological analyses for their assessment, and presents actionable opportunities to push the field forward. By taking a unique, clinical perspective that is absent from technological research on the topic, we find that the state of the art suffers from limited clinical relevance, a lack of overlap between biofluid and physiological approaches, and highly invasive and inaccessible solutions. Future work has the opportunity to maximize the impact of a novel assessment method by incorporating clinical relevance, the holistic nature of micronutrition, and prioritizing accessible and noninvasive systems.
@@
Keywords health sensing  $\cdot$ mobile health  $\cdot$ precision nutrition  $\cdot$ micronutrients  $\cdot$ nutrition assessment  $\cdot$ malnutrition  $\cdot$ point-of-care devices  $\cdot$ accessibility
1 Introduction
@@
Balanced nutrition is important for the development and functioning of the human body and can have many downstream health effects. The World Health Organization (WHO) reports that individuals with proper nutrition have increased lifespans, are more likely to break cycles of poverty, and have lower risks of disease, which impacts productivity and mortality [125]. Malnutrition is the most critical issue related to nutritional intake, with undernutrition associated with about 45% of deaths in children younger than five years old. Anemia is a major resulting disease, affecting 37% of pregnant women and 40% of children under 5. The WHO defines malnutrition as nutrient imbalance, which consists of deficiencies or excesses in essential nutrients [185]. It is a condition that includes several interconnected factors: 1) being overweight, 2) obesity, 3) diet-related noncommunicable diseases like anemia, heart disease, and diabetes, and 4) undernutrition, which encompasses being underweight, wasting, and stunting, as well as a lack of essential micronutrients.
@@
According to Hummell and Cummings [68], malnutrition can be caused by insufficient intake, malabsorption, acute and chronic diseases, increased nutrient need, and weight loss surgery. The impact of acute or chronic diseases and weight loss surgery is self-explanatory, whereas other factors are less visible. Insufficient intake in particular can be influenced by a wide range of factors, making it a challenging problem. These factors include income, access to nutritious foods, culture, and upbringing (malnutrition patterns can be passed down to generations). Insufficient intake is a particularly high risk for low-income families. Furthermore, there is often a lack of awareness about insufficient intake, making behavior change more difficult. Malabsorption can also be a cause, as the absorption of nutrients found in food and supplements is highly individual and can be affected by diseases, genetic issues, and one’s stage of life. In addition, individual nutrient requirements change over time, and several factors can increase nutrient demand, such as growth from infancy to adulthood, pregnancy, lactation, and recovery from illness or other trauma.
@@
Nutrients can be classified into two main categories: macronutrients and micronutrients. Both are essential in precise balance to prevent malnutrition, also referred to as nutritional imbalance. Macronutrients include carbohydrates, fats, and proteins, which are major energy sources and are required in larger amounts than micronutrients [64]. Although required in smaller amounts, optimal intake of micronutrients is crucial. Micronutrients include vitamins and minerals, and they play a critical role in chemical reactions that produce energy from macronutrients acquired from food, as well as other essential functions [50]. Micronutrient intake through diet or supplements is crucial because, unlike macronutrients, our bodies cannot synthesize micronutrients, nor can they be substituted for one another [20]. A precise intake amount is required and slight deviations can result in either deficiency or excess, both with significant health impacts [64]. For these reasons, micronutrient imbalance is a problem that is unique and deserving of attention.
@@
The estimated number of people with micronutrient deficiency is 2 billion worldwide [177]. It is also estimated that micronutrient deficiencies are the cause of between 425,000 and 745,000 deaths in children under five years old [21]. The population of the United States (US) has significant risk of micronutrient deficiency due to the prevalence of a high-energy, low-nutrient diet [50, 42]. The US’ National Health and Nutrition Examination Survey (NHANES) estimates that 31% of the US population is at risk for micronutrient deficiency, with calcium, potassium, iron, and vitamins A, D, C, and E being of particular concern [50, 42]. Since NHANES is a survey of self-reported dietary intake with little biochemical testing, it is likely that this number is actually underestimating the true burden in the US. Deficiencies can cause developmental issues, metabolic disorders, impaired immune system, altered endocrine and cognitive functioning, chronic disease, and more. For example, low magnesium intake is associated with a greater risk of chronic diseases such as cardiovascular disease (CVD) [50]. Those most at risk for micronutrient deficiencies include children less than five years old, pregnant women, victims of chronic disease, and those living in developing nations [42, 23]. Risks of inadequate intake are exacerbated since different micronutrients are needed in different amounts at different stages in life [20]. In the first 1000 days of life, iron, iodine, folate, and vitamin D have high dietary requirements, and a failure to meet them could result in poor physical and cognitive development. During adolescence, iron, calcium, folate, and vitamin D intake is critical, especially for those who menstruate. Pregnancy sees an increase in iron, folate, vitamin B12, and vitamin D requirements. The elderly are more at risk for vitamin D, B12, and B6 deficiency. Medications that influence micronutrient status or absorption must also be more heavily considered during this stage. Micronutrient deficiencies rarely manifest alone, and it is common to see multiple deficiencies arise simultaneously [12].
@@
Like deficiency, micronutrient excess can also lead to health detriments, with the risk increasing as intake levels surpass an individual’s upper limit [56]. Toxicity from micronutrient excess can occur in any person who exceeds this limit. However, there is an exceptional risk for vulnerable populations, namely infants, young children, and pregnant women [70, 135, 56, 46]. For example, an excess of iron increases the risk of diarrhea, sepsis, meningitis, and gut inflammation for infants and young children, with a lethal dose of 150 mg/kg [135]. Similarly, surplus iron can lead to an increased risk of gestational and type 2 diabetes for pregnant women [135]. Excess of vitamins A and D, calcium, and iodine additionally pose elevated health risks for these populations [135, 46]. Individuals in these groups could be affected by micronutrient toxicity because they are more likely to be taking supplements to address an initial deficit [22, 56]. When coupled with population-level nutrition interventions, such as food fortification, one’s micronutrient intake increases, potentially surpassing the recommended value. Research suggests that ingesting an excess of micronutrients through diet alone is unlikely and may only occur for those who take supplements [46, 22, 135]. There is minimal research on micronutrient excess because it is usually limited to the above population groups. As a result, the work covered in this review largely focuses on micronutrient deficiency. For more information on excessive micronutrient intake and its specific effects for each micronutrient, we recommend referring to an up-to-date nutrition overview (e.g. Espinosa-Salas and Gonzalez-Arias [46]).
@@
When a micronutrient imbalance is identified, it is addressed through intervention. This involves modifying diet, providing micronutrient supplements, or on a population scale, fortifying food products such as iodized salt and fortified flour [12]. Interventions need to be carefully planned and monitored to avoid providing too little or too much of a particular micronutrient. There is a strong need for tools that can help guide intervention programs to effectively reach at-risk populations while minimizing the impact on those who have sufficient nutrient levels. This requires employing "different risk assessment methods to make the monitoring process more efficient, reliable, and cost-effective" [12]. Nutritional status assessment is a way to evaluate an individual’s overall nutritional health and is necessary to identify and address any imbalances. This review focuses on micronutrient status assessment methods in particular because we find that existing methods are insufficient and new techniques are under-researched. The visibility of micronutrient status suffers as a result of this lack of attention, further perpetuating associated health issues. An ideal assessment method is accessible and noninvasive. An accessible method is one that is available to more individuals, without compromising clinical effectiveness. This can be achieved through improvements in cost, efficiency, or mobility. Such a method should ultimately alleviate the need for laboratory tests, decrease reliance on high-effort surveys (such as dietary intake logging), and increase efficiency and accuracy for clinicians. A noninvasive method would either eliminate the need for a biofluid sample or obtain it in an unobtrusive manner. An ideal method could obtain micronutrient status regularly, without much effort for the patient. In our review, we aim to consider both clinical and technical aspects of micronutrition assessment and explore the potential for novel, accessible, and noninvasive methods. We argue that such an interdisciplinary approach is key to finding optimal solutions.
1.1 Paper Scope
@@
Few existing papers that review novel micronutrient assessment methods are comprehensive and clinically relevant [172, 156, 121, 78, 67, 25]. Campuzano et al. [25] and Kalita et al. [78] focus particularly on electrochemical sensors, with the latter considering few vitamins and mostly out of body status. Shi et al. [156] concentrate specifically on wearable sensors, and Nimbkar et al. [121] focus on microfluidic assessment. The novelty of our paper lies in its comprehensive review of accessible and noninvasive micronutrient assessment methods, uniquely emphasizing clinical relevance—an aspect often overlooked in prior work. It includes a background on micronutrients, tailored for a non-clinical perspective. We also explore a wide range of techniques for assessing physiology and biofluids, including assay-based technologies, electrochemistry-based methods, spectroscopy-based approaches, and analytic methods utilizing artificial intelligence (AI) and machine learning (ML). This review contributes several relevant tables throughout the paper and in the Appendix providing a comprehensive reference about micronutrients and their assessment. Ultimately, our paper calls for actionable potential opportunities to advance these methods for clinical use.
@@
With the aim of pushing innovative research towards an accessible, noninvasive method of assessing micronutrient status, this review consists mostly of methods that are both accessible and noninvasive, or at least one of these. We additionally include few methods that are neither accessible nor noninvasive, yet are interesting and valuable for future work within this scope. This review does not include methods for assessing micronutrient levels in entities outside of humans, such as food or pharmaceuticals. Our focus is on vitamins and trace minerals (micronutrients), thus we do not discuss nutrients such as proteins, fats, and carbohydrates (macronutrients) or major minerals (e.g. potassium, calcium, etc.). These are sometimes considered separate from micronutrients because of their higher intake requirements and larger quantities in the body.
@@
This review is structured as follows. We first present background information in Section 2. Next, we discuss micronutrient status assessment methods based on biofluid analysis in Section 3, and physiology-based methods in Section 4. We present suggestions for future work in Section 5. Finally, we end with a conclusion in Section 6, followed by reference tables in Section A.
2 Background
In this section we describe micronutrient characteristics, their presence in biofluids, and related physiological effects. We begin with a summary of micronutrients, their characteristics, their presence in biofluids, and their impacts on our physiological functioning. Such a background is necessary to contextualize solutions to micronutrient status assessment because of the variety of possible approaches.
2.1 Micronutrient Characteristics

### v2

@@ Abstract.
Nutrients are critical to the functioning of the human body and their imbalance can result in detrimental health concerns. The majority of nutritional literature focuses on macronutrients, often ignoring the more critical nuances of micronutrient balance, which require more precise regulation. Currently, micronutrient status is routinely assessed via complex methods that are arduous for both the patient and the clinician. To address the global burden of micronutrient imbalance, innovations in assessment must be accessible and noninvasive. In support of this task, this article synthesizes useful background information on micronutrients themselves, reviews the state of biofluid and physiological analyses for their assessment, and presents actionable opportunities to push the field forward. By taking a unique, clinical perspective that is absent from technological research on the topic, we find that the state of the art suffers from limited clinical relevance, a lack of overlap between biofluid and physiological approaches, and highly invasive and inaccessible solutions. We present opportunities for future work to maximize the impact of a novel assessment method by incorporating clinical relevance, the holistic nature of micronutrition, and prioritizing accessible and noninvasive systems.
@@
health sensing, mobile health, precision nutrition, micronutrients, nutrition assessment, malnutrition, point-of-care devices, accessibility
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXX
1. Introduction
@@
Balanced nutrition is important for the development and functioning of the human body and can have many downstream health effects. The World Health Organization (WHO, Table 12) reports that individuals with proper nutrition have increased lifespans, are more likely to break cycles of poverty, and have lower risks of disease, which impacts productivity and mortality (Organization, 2024). Malnutrition is the most critical issue related to nutritional intake, with undernutrition associated with about 45% of deaths in children younger than five years old. Anemia is a major resulting disease, affecting 37% of pregnant people and 40% of children under 5. The WHO defines malnutrition as nutrient imbalance, which consists of deficiencies or excesses in essential nutrients (World Health Organization, 2024). It is a condition that includes several interconnected factors: 1) being overweight, 2) obesity, 3) diet-related noncommunicable diseases like anemia, heart disease, and diabetes, and 4) undernutrition, which encompasses being underweight, wasting, and stunting, as well as a lack of essential micronutrients. Micronutrients are the vitamins (e.g. vitamins A, C, B12) and minerals (e.g. iron, calcium, iodine) which are critical to everyday human bodily function and can only be provided via dietary intake (i.e. they are not produced by the body).
@@
According to Hummell and Cummings (2022), malnutrition can be caused by insufficient intake, malabsorption, acute and chronic diseases, increased nutrient need, and weight loss surgery. The impact of acute or chronic diseases and weight loss surgery is self-explanatory, whereas other factors are less visible. Insufficient intake in particular can be influenced by a wide range of variables, making it a challenging problem. These factors include income, access to nutritious foods, culture, and upbringing (malnutrition patterns can be passed down to generations). Insufficient intake is a particularly high risk for low-income families. Furthermore, there is often a lack of awareness about insufficient intake, making behavior change more difficult. Malabsorption can also be a cause, as the absorption of nutrients found in food and supplements is highly individual and can be affected by diseases, genetic issues, and one’s stage of life. In addition, individual nutrient requirements change over time, and several factors can increase nutrient demand, such as growth from infancy to adulthood, pregnancy, lactation, and recovery from illness or other trauma.
@@
Nutrients can be classified into two main categories: macronutrients and micronutrients. Both are essential in precise balance to prevent malnutrition, also referred to as nutritional imbalance. Macronutrients include carbohydrates, fats, and proteins, which are major energy sources and are required in larger amounts than micronutrients (Heitman, 2022). Although required in smaller amounts, optimal intake of micronutrients is crucial. Micronutrients include vitamins and minerals, and they play a critical role in chemical reactions that produce energy from macronutrients acquired from food, as well as other essential bodily functions (Organization, 2025). Micronutrient intake through diet or supplements is crucial because our bodies cannot synthesize micronutrients, nor can they be substituted for one another (Biesalski Hans and Jana, 2018). A precise intake amount is required and slight deviations can result in either deficiency or excess, both with significant health impacts (Heitman, 2022). For these reasons, micronutrient imbalance is a problem that is unique and deserving of attention.
@@
The estimated number of people with micronutrient deficiency is 2 billion worldwide (Von Grebmer et al., 2014). It is estimated that micronutrient deficiencies are the cause of between 425,000 and 745,000 deaths in children under five years old (Brown et al., 2021). An estimated 56% of preschool-aged children and 69% of non-pregnant females of reproductive age worldwide suffer from at least one micronutrient deficiency (Stevens et al., 2022). Globally, the most widespread micronutrient deficiencies occur in iron, iodine, folate, vitamin A, and zinc (Bailey et al., 2015). The population of the United States (US) has significant risk of micronutrient deficiency due to the prevalence of a high-energy, low-nutrient diet (for Functional Medicine, 2020; Drake, 2018). The US’ National Health and Nutrition Examination Survey (NHANES) estimates that 31% of the US population is at risk for micronutrient deficiency, with calcium, potassium, iron, and vitamins A, D, C, and E being of particular concern (for Functional Medicine, 2020; Drake, 2018). Since NHANES is a survey of self-reported dietary intake with minimal biochemical testing, it is likely that this number is actually underestimating the true burden in the US (Drake, 2018).
Deficiencies can cause developmental issues, metabolic disorders, impaired immune system, altered endocrine and cognitive functioning, chronic disease, and more. For example, high magnesium intake and circulating status (via urine) is directly associated with a greater risk of cardiovascular disease (CVD) (Rosique-Esteban et al., 2018). While developed countries are certainly at a risk for micronutrient deficiencies, those most at risk are people living in underdeveloped countries. Populations that also bear a significant burden include developing countries as well as children less than five years old, pregnant people, and victims of chronic disease (Drake, 2018; Bruins et al., 2016; Stevens et al., 2022; Organization, 2025). Risks of inadequate intake are exacerbated since different micronutrients are needed in different amounts at different stages in life (Biesalski Hans and Jana, 2018). Vitamin A, iron, and zinc are among the most significant micronutrient deficiencies observed in children (Stevens et al., 2022). In the first 1000 days of life, iron, iodine, folate, and vitamin D have high dietary requirements, and a failure to meet them could result in poor physical and cognitive development. During adolescence, iron, calcium, folate, and vitamin D intake is critical, especially for those who menstruate. Pregnancy sees an increase in iron, folate, vitamin B12, and vitamin D requirements. The elderly are more at risk for vitamin D, B12, and B6 deficiency, and medications that influence micronutrient status or absorption must be heavily considered during this stage. Micronutrient deficiencies rarely manifest alone, and it is common to see multiple deficiencies arise simultaneously (Bailey et al., 2015).
@@
Like deficiency, micronutrient excess can also lead to health detriments, with the risk increasing as intake levels surpass an individual’s upper limit (Gernand, 2019). Toxicity from micronutrient excess can occur in any person who exceeds this limit. However, there is an exceptional risk for vulnerable populations, namely infants, young children, and pregnant people (Ishaq et al., 2024; Pike and Zlotkin, 2019; Gernand, 2019; Espinosa-Salas and Gonzalez-Arias, 2024). For example, an excess of iron increases the risk of diarrhea, sepsis, meningitis, and gut inflammation for infants and young children, with a lethal dose of 150 mg/kg (Pike and Zlotkin, 2019). Similarly, surplus iron can lead to an increased risk of gestational and type 2 diabetes for pregnant people (Pike and Zlotkin, 2019). Elevated health risks are also a concern with an excess of vitamin D in infants, calcium in pregnant people, vitamin A in infants, young children, and pregnant people, and iodine in all three of these groups in addition to breastfeeding parents (Pike and Zlotkin, 2019; Espinosa-Salas and Gonzalez-Arias, 2024). Individuals who are more likely to be taking supplements (especially multivitamins) to address an initial deficit could also be at a greater risk for micronutrient toxicity (Bruins et al., 2015; Gernand, 2019). Research suggests that ingesting an excess of micronutrients through diet alone is unlikely and may only occur for those who take supplements (Espinosa-Salas and Gonzalez-Arias, 2024; Bruins et al., 2015; Pike and Zlotkin, 2019). An excess of certain micronutrients can interfere with the absorption of others, potentially leading to secondary deficiencies. For instance, elevated calcium levels may reduce the absorption of magnesium. Additional information on micronutrient interactions are provided in Tables LABEL:tab:_char-ws, LABEL:tab:_char-fs, and LABEL:tab:_char-min. There is minimal research on micronutrient excess because it is usually limited to the above population groups. As a result, the work covered in this review largely focuses on micronutrient deficiency. For more information on excessive micronutrient intake and its specific effects for each micronutrient, we recommend referring to an up-to-date nutrition overview (e.g. Espinosa-Salas and Gonzalez-Arias (2024)).
@@
When a micronutrient imbalance is identified, it is addressed through intervention. This involves modifying diet, providing micronutrient supplements, or on a population scale, fortifying food products such as iodized salt and fortified flour (Bailey et al., 2015). Interventions need to be carefully planned and monitored to avoid providing too little or too much of a particular micronutrient. There is a strong need for tools that can help guide intervention programs to effectively reach at-risk populations while minimizing the impact on those who have sufficient nutrient levels. This requires employing ”different risk assessment methods to make the monitoring process more efficient, reliable, and cost-effective” (Bailey et al., 2015). Nutritional status assessment is a way to evaluate an individual’s overall nutritional health and is necessary to identify, prevent, and address any imbalances. This review focuses on micronutrient status assessment methods in particular because we find that existing methods are insufficient and new techniques are under-researched.
We propose that an ideal assessment method is both accessible and noninvasive. An accessible method is one that is available to more individuals, while maintaining clinical effectiveness. This can be achieved through improvements in cost, efficiency, or mobility. Such a method should ultimately alleviate the need for laboratory tests, decrease reliance on high-effort and subjective surveys (such as dietary intake logging), and increase efficiency and accuracy for clinicians. A noninvasive method would either eliminate the need for a biofluid sample or obtain it in an unobtrusive manner. Noninvasiveness is therefore interconnected with accessibility, as a method that is invasive (e.g. requiring a blood draw) is less accessible (e.g. places burden on the patient and on the clinician). An ideal method that is both accessible and noninvasive could obtain micronutrient status regularly, without much effort for the patient or clinician. To maintain effective and clinically-relevant assessment methods, it is also essential to balance the accessibility and noninvasiveness of a method with its ability to be a sensitive (high true positive rate) and specific (high true negative rate) assessment of micronutrient status. Because every country in the world is impacted by malnutrition (World Health Organization, 2024), it is clear that an accessible and noninvasive assessment method is needed to enable routine assessment globally. This is especially true in countries with high-fat, low-nutrient diets such as the US, in low- and middle-income nations, and in areas of the world where complex, blood-based analyses are generally infeasible. In our review, we aim to consider both clinical and technical aspects of micronutrition assessment and explore the potential for novel, accessible, and noninvasive methods. We argue that such an interdisciplinary approach is key to finding optimal solutions.
In the subsequent sections, we discuss biofluid and physiological analysis methods for micronutrient status assessment. Within the context of this paper, biofluid analysis refers to techniques that measure the amount of a micronutrient biomarker present in a biofluid as a means to infer the underlying nutritional status of an individual, whereas physiological analysis focuses on the bodily signs and signals which are symptomatic a micronutritional imbalance. For the latter, we focus on providing a brief overview of clinical physical assessment in nutrition and describe relevant applications of optical sensing.
1.1. Paper Scope
Few existing papers that review novel micronutrient assessment methods are comprehensive and clinically relevant (Udhani et al., 2023; Shi et al., 2022; Nimbkar et al., 2024; Kalita et al., 2022; Huey et al., 2022; Campuzano et al., 2024). Campuzano et al. (2024) and Kalita et al. (2022) focus particularly on electrochemical sensors, with the latter considering few vitamins and mostly out-of-body status. Huey et al. (2022) focus on vitamin A assessment and emphasizes the limitations of relying on centralized laboratories and specialized equipment, underscoring the need for more portable and accessible diagnostic methods. Nimbkar et al. (2024) focus on microfluidic assessment and Shi et al. (2022) concentrate specifically on wearable sensors. Udhani et al. (2023) provide a detailed review on biomarker-based micronutrient detection, focusing on the analytical chemistry aspect.
@@ 1.1. Paper Scope
The novelty of our paper lies in its comprehensive review of accessible and noninvasive micronutrient assessment methods, uniquely emphasizing clinical relevance—an aspect often overlooked in prior work. It includes a background on micronutrients, tailored for a non-clinical perspective. We also explore a wide range of techniques for assessing biofluids and physiology, including assay-based technologies, electrochemistry-based methods, spectroscopy-based approaches, and analytic methods utilizing machine learning (ML). This review offers a high-level summary, with detailed methods available in dedicated sources. We contribute several relevant tables throughout the paper and provide a comprehensive reference about micronutrients and their assessment in the Appendix. Ultimately, this review calls for actionable potential opportunities to advance these methods.
@@
This review consists mostly of methods that are both accessible and noninvasive, or at least one of these. We additionally include few methods that are neither accessible nor noninvasive, yet are interesting and valuable for future work within this scope. This review does not include methods for assessing micronutrient levels in entities outside of humans, such as food or pharmaceuticals. We note that while hair and nail samples can be valid subjects of micronutrient analysis, we focus on biofluids in our review of assessment techniques since nutritional analyses of hair and nails measure the metabolism of a nutrient in cells rather than the in-body status of the nutrient (Han et al., 2016; Bali et al., 2023). Our focus is on vitamins and trace minerals (micronutrients), thus we do not discuss nutrients such as proteins, fats, and carbohydrates (macronutrients) or major minerals (e.g. potassium, calcium, etc.). These are sometimes considered separate from micronutrients because of their higher intake requirements and larger quantities within the body. Lastly, this review does not include alternative medicine techniques or methods, as this is out of the scope.
@@
This review is structured as follows. We first present background information in Section 2. Next, we discuss micronutrient status assessment methods based on biofluid analysis in Section 3, and briefly touch on physiology-based methods in Section 4, focusing on optical sensing techniques. We summarize the current gaps in the literature and present suggestions for future work in Section 5. Finally, we end with a conclusion in Section 6, followed by reference tables in the Appendix (A).
2. Background
In this section we describe micronutrient characteristics, their presence in biofluids, and their impacts on physiological functioning. Such a background is necessary to contextualize solutions to emerging micronutrient status assessment methods because of the variety of possible approaches. In-body assessment methods refer to the levels of a micronutrient present within the body, reflecting the current internal state of an individual’s micronutrition. On the other hand, out-of-body assessment methods involve external indicators used to infer internal status, such as physiological symptoms measurable through wearable sensors or physical exams. We focus on both approaches, as each provides relevant insights into micronutrient status and offers the potential for non-invasive and accessible assessment methods.
2.1. Micronutrient Characteristics